# Towards the Colmez Conjecture

### v1

@@ 2. Partial CM Types and the Colmez Conjecture
Let $E$ be a CM-field, which is a totally imaginary quadratic extension of a totally real number field, and let its degree be $[E:\mathbb{Q}]=2g$ and its ring of integers be $\mathcal{O}_{E}$. There exists a unique complex conjugation automorphism of $E$, which we denote $c$, and for $\sigma\in H_{E}$, let $\overline{\sigma}\coloneqq\sigma\circ c$. A CM-type $\Phi\subset H_{E}$ is a subset such that $\Phi\sqcup\overline{\Phi}=H_{E}$. Note that giving a CM-type $\Phi$ is the same as giving an isomorphism $E\otimes\mathbb{R}\to\mathbb{C}^{g}$.
@@
Let $A$ be an abelian variety with CM of type $(\mathcal{O}_{E},\Phi)$ and let $\tau\in\Phi$. Colmez uses the pairing between $H_{\operatorname{dR}}^{1}(A/\mathbb{C})$ and $H_{1}(A(\mathbb{C}),\mathbb{Z})$ to define a $\tau$-component of the Faltings height, which we calls $ht(\Phi,\tau)$. He then proves that $ht(\Phi,\tau)$ can be extended to a linear height function on $\mathcal{CM}^{0}$.
@@
$b_{\tau,\rho}(g)=\begin{cases}\frac{1}{2}&\text{if $g\tau=\rho$,}\\ -\frac{1}{2}&\text{if $g\tau=\overline{\rho}$,}\\ 0&\text{otherwise.}\end{cases}$
We note that we can write this height explicitly in terms of Faltings heights of CM abelian varieites.
@@
Let $h(\Phi)$ denote the Faltings height of an abelian variety with CM of type $(\mathcal{O}_{E},\Phi)$. Let $[E:\mathbb{Q}]=2g$. Then
@@
The second summation is 00 because $b_{\tau,\rho}+b_{\tau,\overline{\rho}}=0$ for any $\tau,\rho\in H_{E}$. Meanwhile, noting that $A_{\Phi}=g/2+\sum_{\tau,\rho\in\Phi}b_{\tau,\rho}$, we can simplify the right hand side as
$\frac{1}{2^{g-\lvert\phi\rvert}}\sum_{\Phi\supset\phi}A_{\Phi}-\frac{g-\lvert\phi\rvert}{g2^{g}}\sum_{\Phi^{\prime}}A_{\Phi^{\prime}}=\frac{\lvert\phi\rvert}{2}+\sum_{\tau,\rho\in\phi}b_{\tau,\rho}+\frac{1}{2}\sum_{\tau|_{F}\not\in\Sigma}b_{\tau,\tau}-\frac{g-\lvert\phi\rvert}{2g}\sum_{\tau\in H_{E}}b_{\tau,\tau}.$
However, it is clear that $b_{\tau,\tau}$ is independent of $\tau$ (and is equal to $\frac{1}{2g}\operatorname{Ind}_{E/F}^{E/\mathbb{Q}}\chi_{E/F}$), and thus the last two summations cancel giving the desired equality. ∎
Moreover, using the comparison between Faltings height and Colmez’s height function, we can express the Faltings height of a CM abelian varieties in terms of partial CM-types of order $2$. Note that this does cannot simply be determined from the Proposition given above.
@@
To prove Conjecture 1.2, it suffices to prove the result for all pairs
@@
In [YZ18, §2], they also define the discriminant of a CM-type, which we temporarily denote $d_{\Phi,YZ}$. Our discriminant is gotten by normalizing so $d_{\Phi}=d_{\Phi,YZ}^{1/[E_{\Phi}:\mathbb{Q}]}$.
We will also use the following Proposition of Colmez.
@@
Let $E/\mathbb{Q}_{p}$ be a finite extension and $L\subset\overline{\mathbb{Q}_{p}}$ be a finite Galois extension of $\mathbb{Q}_{p}$ that contains all the conjugates of $E$. Let $f_{E}\coloneqq[\mathcal{O}_{E}/\mathfrak{m}_{E}:\mathbb{F}_{p}]$ and let $\varphi_{p}$ denote the Frobenius map. Fix an embedding $E\subset\overline{\mathbb{Q}_{p}}$ and for $\sigma\in H_{E}$, let $i(\sigma)\in\mathbb{Z}/f_{E}\mathbb{Z}$ be such that $\sigma$ induces the map $\varphi_{p}^{i(\sigma)}$ on the residue field. Let $\sigma,\tau\in H_{E}$ and let $a_{\sigma,\tau}\colon\operatorname{Gal}(\overline{\mathbb{Q}_{p}}/\mathbb{Q}_{p})\to\mathbb{Q}$ be the function such that $a_{\sigma,\tau}(g)=1$ if $g\sigma=\tau$ and 00 otherwise.
@@
$\mu_{Art}(A_{\phi}^{0})=\frac{\lvert\phi\rvert}{2[E:\mathbb{Q}]}\log d_{E}-\frac{1}{2}\log d_{\phi}+\frac{1}{4}\log d_{\Sigma}.$
@@
$A_{\phi}(g)=\lvert\phi\cap g\phi\rvert+\frac{\lvert\phi\rvert}{2}-\frac{1}{2}\lvert\Sigma\cap g\Sigma\rvert,$
which written in terms of $a_{\sigma,\tau}$ is
@@
which recovers the log terms of [AGHMP18, YZ18].
@@
Assume Conjecture 1.2. Let $\sigma,\tau\in H_{E}$ be two distinct places such that $\sigma\neq\overline{\tau}$ and let $\phi=\{\sigma,\tau\}$. Let $L\subset\mathbb{Q}^{\mathrm{cm}}$ be a finite Galois CM-extension of $\mathbb{Q}$ that contains all the conjugates of $E$ and let $H\coloneqq\operatorname{Gal}(L/\sigma(E))\subset\operatorname{Gal}(L/\mathbb{Q})$ be the stabilizer of $\sigma(E)\subset L$. Let $g_{0}\in\operatorname{Gal}(L/\mathbb{Q})$ be an element such that $g\sigma=\tau$ and let $c\in\operatorname{Gal}(L/\mathbb{Q})$ denote complex conjugation. Then
$h(\phi)=$ $-\frac{1}{g}\frac{L^{\prime}(\chi_{E/F},0)}{L(\chi_{E/F},0)}-\frac{1}{2}\sum_{\chi}\left(\sum_{h\in H}\chi(g_{0}h)+\chi(g_{0}ch)+\frac{1}{\chi(g_{0}h)}+\frac{1}{\chi(g_{0}ch)}\right)\frac{L^{\prime}(\chi,0)}{L(\chi,0)}$
$-\frac{1}{4[E:\mathbb{Q}]}\log d_{E}+\frac{1}{4}\log d_{\phi}-\frac{1}{8}\log d_{\Sigma}+\log 2\pi,$
@@
The discriminant term is calculated in Proposition 2.9 and so it suffices to express $A_{\phi}^{0}$ in terms of Artin characters in order calculate what the $L$-function terms looks like. We have the decomposition
@@
$\langle\chi,b_{\sigma,\tau}^{0}\rangle=\sum_{h\in H}\chi(g_{0}h).$
@@
$b_{\sigma,\tau}=\frac{1}{2}\mathbbm{1}_{g_{0}H}-\frac{1}{2}\mathbbm{1}_{g_{0}cH},$
@@
$\langle\chi,\mathbbm{1}_{g}^{0}\rangle=\chi(g)$
for any $g\in\operatorname{Gal}(L/\mathbb{Q})$. ∎
@@ 3.1. Faltings Height
Let $A$ be an abelian variety of dimension $g$ defined over a number field $K\subset\mathbb{C}$ with semi-stable reduction (enlarging the field $K$ if necessary) and let $\mathcal{A}$ be the Néron model over $\mathcal{O}_{K}$ and let $\Omega_{\mathcal{A}/\mathcal{O}_{K}}$ denote the sheaf of relative differentials. Let $\Omega(\mathcal{A})\coloneqq H^{0}(\mathcal{A},\Omega_{\mathcal{A}/\mathcal{O}_{K}})$ and let $\omega(\mathcal{A})\coloneqq H^{0}(\mathcal{A},\Omega_{\mathcal{A}/\mathcal{O}_{K}}^{g})$ denote the Hodge bundle of $\mathcal{A}$. This is rank $1$ projective $\mathcal{O}_{K}$-module and $\omega(A)\coloneqq\omega(\mathcal{A})\otimes_{\mathcal{O}_{K}}K\cong H^{0}(A,\Omega_{A/K}^{g})$ is a $1$-dimensional $K$-vector space. For each $\omega\in\omega(\mathcal{A})$ and $\sigma\in H_{K}$, we let $\omega^{\sigma}$ be the base change of $\omega$ viewed as an element of $\omega(\mathcal{A})\otimes_{\mathcal{O}_{K},\sigma}R\cong H^{0}(A^{\sigma},\Omega^{g}_{A^{\sigma}/R})$, where $R=\overline{\mathbb{Q}_{p}}$ or $\mathbb{C}$ and $A^{\sigma}\coloneqq A\otimes_{K,\sigma}R$. We view $\omega(\mathcal{A})$ as a lattice in $H^{0}(A,\Omega_{A/K}^{g})$. This lattice allows us to define a valuation on $H^{0}(A,\Omega_{A/K}^{g})$ for each rational prime $p$ and embedding $\sigma\in\operatorname{Hom}(K,\overline{\mathbb{Q}_{p}})$ by setting $v_{p}(\omega^{\sigma})=0$ if $\omega^{\sigma}$ generates $\omega(\mathcal{A})\otimes_{\mathcal{O}_{K},\sigma}\mathcal{O}_{p}$ as an $\mathcal{O}_{p}$-module and $v_{p}(\alpha\omega^{\sigma})=v_{p}(\sigma(\alpha))+v_{p}(\omega^{\sigma})$ for any $\alpha\in K$. This gives us the local contribution to the height.
We now define a norm for each $\sigma\colon K\to\mathbb{C}$. We can view $\omega^{\sigma}\in H^{0}(A^{\sigma},\Omega_{A^{\sigma}/\mathbb{C}}^{g})$ as a holomorphic $g$-form on $A^{\sigma}$ and we set
@@
for a choice of $\omega\in\omega(A)\backslash\{0\}$. This is well defined and independent of the choice of $\omega$ by the product formula.

### v2

@@ 2. Partial CM Types and the Colmez Conjecture
Let $E$ be a CM-field, which is a totally imaginary quadratic extension of a totally real number field, and let its degree be $E:\mathbb{Q}]=2g$ and its ring of integers be $\mathcal{O}_{E}$. There exists a unique complex conjugation automorphism of $E$, which we denote $c$, and for $\sigma\in H_{E}$, let $\overline{\sigma}\coloneqq\sigma\circ c$. A CM-type $\Phi\subset H_{E}$ is a subset such that $\Phi\sqcup\overline{\Phi}=H_{E}$. Note that giving a CM-type $\Phi$ is the same as giving an isomorphism $E\otimes\mathbb{R}\to\mathbb{C}^{g}$.
@@
Let $A$ be an abelian variety with CM of type $(\mathcal{O}_{E},\Phi)$ and let $\tau\in\Phi$. Colmez uses the pairing between $H_{\operatorname{dR}}^{1}(A/\mathbb{C})$ and $H_{1}(A(\mathbb{C}),\mathbb{Z})$ to define a $\tau$-component of the Faltings height, which we denote $ht(\Phi,\tau)$. He then proves that $ht(\Phi,\tau)$ can be extended to a linear height function on $\mathcal{CM}^{0}$.
@@
$b_{\tau,\rho}(x)=\begin{cases}\frac{1}{2}&\text{if $x\tau=\rho$,}\\ -\frac{1}{2}&\text{if $x\tau=\overline{\rho}$,}\\ 0&\text{otherwise.}\end{cases}$
We note that we can write this height explicitly in terms of Faltings heights of CM abelian varieties.
@@
Let $h(\Phi)$ denote the Faltings height of an abelian variety with CM of type $(\mathcal{O}_{E},\Phi)$. Let $E:\mathbb{Q}]=2g$. Then
@@
The second summation is 0 because $b_{\tau,\rho}+b_{\tau,\overline{\rho}}=0$ for any $\tau,\rho\in H_{E}$. Meanwhile, noting that $A_{\Phi}=g/2+\sum_{\tau,\rho\in\Phi}b_{\tau,\rho}$, we can simplify the right hand side as
$\frac{1}{2^{g-\lvert\phi\rvert}}\sum_{\Phi\supset\phi}A_{\Phi}-\frac{g-\lvert\phi\rvert}{g2^{g}}\sum_{\Phi^{\prime}}A_{\Phi^{\prime}}=$ $\frac{\lvert\phi\rvert}{2}+\sum_{\tau,\rho\in\phi}b_{\tau,\rho}$
$+\frac{1}{2}\sum_{\tau|_{F}\not\in\Sigma}b_{\tau,\tau}-\frac{g-\lvert\phi\rvert}{2g}\sum_{\tau\in H_{E}}b_{\tau,\tau}.$
It is clear that $b_{\tau,\tau}$ is independent of $\tau$ (and moreover is equal to $\frac{1}{2g}\operatorname{Ind}_{E/F}^{E/\mathbb{Q}}\chi_{E/F}$), and thus the last two summations cancel giving the desired equality. ∎
Moreover, using the comparison between the Faltings height and Colmez’s height function, we can express the Faltings height of a CM abelian variety in terms of partial CM-types of order $2$. Note that this cannot be directly deduced from the proposition above.
@@
To prove Conjecture 1.3, it suffices to prove the result for all pairs of distinct embeddings $\rho,\tau\in H_{E}$. Namely, to prove that
@@
In [YZ18, §2], they also define the discriminant of a CM-type, which we temporarily denote $d_{\Phi,YZ}$. We normalize our discriminant so $d_{\Phi}=d_{\Phi,YZ}^{1/[E_{\Phi}:\mathbb{Q}]}$.
Recall that $\mu_{Art}\in\mathcal{CM}^{0}$ was the function given on Artin characters $\chi$ by $\mu_{Art}(\chi)=\log f_{\chi}$. For each finite prime $p\in\mathbb{Z}$, we define the function $\mu_{Art,p}\in\mathcal{CM}^{0}$ as
$\mu_{Art,p}(\chi)=v_{p}(f_{\chi}),$
where $v_{p}$ is the $p$-adic valuation normalized so that $v_{p}(p)=1$. In this way, we have
$\mu_{Art}(a)=\sum_{p}\mu_{Art,p}(a)\log p.$
There is the following proposition of Colmez.
@@
Let $E/\mathbb{Q}_{p}$ be a finite extension and $L\subset\overline{\mathbb{Q}_{p}}$ be a finite Galois extension of $\mathbb{Q}_{p}$ that contains all the conjugates of $E$. Let $f_{E}\coloneqq[\mathcal{O}_{E}/\mathfrak{m}_{E}:\mathbb{F}_{p}]$ and let $\varphi_{p}$ denote the Frobenius map. Fix an embedding $E\subset\overline{\mathbb{Q}_{p}}$ and for $\sigma\in H_{E}$, let $i(\sigma)\in\mathbb{Z}/f_{E}\mathbb{Z}$ be such that $\sigma$ induces the map $\varphi_{p}^{i(\sigma)}$ on the residue field. Let $\sigma,\tau\in H_{E}$ and let $a_{\sigma,\tau}\colon\operatorname{Gal}(\overline{\mathbb{Q}_{p}}/\mathbb{Q}_{p})\to\mathbb{Q}$ be the function such that $a_{\sigma,\tau}(x)=1$ if $x\sigma=\tau$ and 0 otherwise.
@@
Let $d_{E}\in\mathbb{Z}$ denote the absolute discriminant of $E$. Then
$\mu_{Art}(A_{\phi}^{0})=\frac{\lvert\phi\rvert}{2[E:\mathbb{Q}]}\log|d_{E}|-\frac{1}{2}\log d_{\phi}+\frac{1}{4}\log d_{\Sigma}.$
@@
$A_{\phi}(x)=\lvert\phi\cap x\phi\rvert+\frac{\lvert\phi\rvert}{2}-\frac{1}{2}\lvert\Sigma\cap x\Sigma\rvert$,
which, written in terms of $a_{\sigma,\tau}$, is
@@
which recovers the log terms of [AGH+18, YZ18].
@@
Assume Conjecture 1.3. Let $\sigma,\tau\in H_{E}$ be two distinct places such that $\sigma\neq\overline{\tau}$ and let $\phi=\{\sigma,\tau\}$. Let $L\subset\mathbb{Q}^{\mathrm{cm}}$ be a finite Galois CM-extension of $\mathbb{Q}$ that contains all the conjugates of $E$ and let $H\coloneqq\operatorname{Gal}(L/\sigma(E))\subset\operatorname{Gal}(L/\mathbb{Q})$ be the stabilizer of $\sigma(E)\subset L$. Let $x_{\sigma,\tau}\in\operatorname{Gal}(L/\mathbb{Q})$ be an element such that $x_{\sigma,\tau}\sigma=\tau$ and let $c\in\operatorname{Gal}(L/\mathbb{Q})$ denote complex conjugation. Then
$h(\phi)=$ $-\frac{1}{2}\sum_{\chi}\left(\sum_{y\in H}\chi(x_{\sigma,\tau}y)+\chi(x_{\sigma,\tau}cy)+\frac{1}{\chi(x_{\sigma,\tau}y)}+\frac{1}{\chi(x_{\sigma,\tau}cy)}\right)\frac{L^{\prime}(\chi,0)}{L(\chi,0)}$
@@
The discriminant term is calculated in Proposition 2.9 and so it suffices to express $A_{\phi}^{0}$ in terms of Artin characters in order to calculate what the $L$-function terms looks like. We have the decomposition
@@
$\langle\chi,b_{\sigma,\tau}^{0}\rangle=\sum_{h\in H}\chi(x_{\sigma,\tau}y).$
@@
$b_{\sigma,\tau}=\frac{1}{2}\mathbbm{1}_{x_{\sigma,\tau}H}-\frac{1}{2}\mathbbm{1}_{x_{\sigma,\tau}cH},$
@@
$\langle\chi,\mathbbm{1}_{x}^{0}\rangle=\chi(x)$
for any $x\in\operatorname{Gal}(L/\mathbb{Q})$. ∎
Example 2.11.
We note that if $\phi$ is a full CM-type of $E$, then $h(\phi)$ corresponds to the Faltings height of any CM abelian variety of type $(\mathcal{O}_{E},\phi)$. These heights can drastically vary for a fixed $E$, even in the case of $\lvert\phi\rvert=2$, as we now demonstrate.
Let $E=\mathbb{Q}(\sqrt{-1},\sqrt{p})\subset\mathbb{C}$ be the biquadratic field with the embeddings given by $H_{E}=\{1,c,\sigma,c\sigma\}$, where $c$ is complex conjugation on $\mathbb{C}$ and $\sigma$ fixes $i$ but sends $\sqrt{p}\to-\sqrt{p}$. Then there are two non-equivalent CM-types on $E$ given by $\Phi_{1}=\{1,\sigma\}$ and $\Phi_{2}=\{1,\sigma c\}$. However, these are not primitive CM-types and induced from CM-types on $E_{1}=\mathbb{Q}(\sqrt{-1})$ and $E_{2}=\mathbb{Q}(\sqrt{-p})$, respectively. For $i\in\{1,2\}$, let $h(E_{i})$ denote the Faltings height of an elliptic curve with CM by $\mathcal{O}_{E_{i}}$. Then, we have that $h(\Phi_{1})=2h(E_{1})$ and $h(\Phi_{2})=2h(E_{2})$. The former is constant, whereas conjecturally under GRH, the latter grows logarithmically with $p$. We briefly explain. Conjecture 1.3 holds for elliptic curves (as it is just a reformulation of the Chowla–Selberg Formula) and hence we can express the Faltings height $h(E_{2})$ in terms of the logarithmic derivative $\frac{L^{\prime}(\chi_{E_{2}},0)}{L(\chi_{E_{2}},0)}$, where $\chi_{E_{2}}$ is the nontrivial quadratic character associated with $E_{2}$. Taking the logarithmic derivative of the functional equation of the completed $L$-function gives the equation
$h(E_{2})=\frac{1}{4}\log d_{E_{2}}+\frac{1}{2}\frac{L^{\prime}(\chi_{E_{2}},1)}{L(\chi_{E_{2}},1)}-\frac{\gamma}{2}-\log 2\pi,$
where $\gamma$ is Euler’s constant. Under the assumption of GRH, the logarithmic derivative at $1$ term is $O(\log\log d_{E_{2}})$ (see [IMS09, Thm. 3]) and hence the dominant term is $\frac{1}{4}\log d_{E_{2}}$. Thus, for fixed degree $2g=[E:\mathbb{Q}]$, the ratio and difference between two heights of partial CM-types of a given field can be arbitrarily large.
@@ 3.1. Faltings Height
Let $A$ be an abelian variety of dimension $g$ defined over a number field $K\subset\mathbb{C}$, with semi-stable reduction (after possibly enlarging $K$ if necessary). Let $\mathcal{A}$ be the Néron model over $\mathcal{O}_{K}$ and let $\Omega_{\mathcal{A}/\mathcal{O}_{K}}$ denote the sheaf of relative differentials. Let $\Omega(\mathcal{A})\coloneqq H^{0}(\mathcal{A},\Omega_{\mathcal{A}/\mathcal{O}_{K}})$ be the $\mathcal{O}_{K}$-module and let $\omega(\mathcal{A})\coloneqq H^{0}(\mathcal{A},\Omega_{\mathcal{A}/\mathcal{O}_{K}}^{g})$ denote the Hodge bundle of $\mathcal{A}$. This is rank $1$ projective $\mathcal{O}_{K}$-module and $\omega(A)\coloneqq\omega(\mathcal{A})\otimes_{\mathcal{O}_{K}}K\cong H^{0}(A,\Omega_{A/K}^{g})$ is a $1$-dimensional $K$-vector space. For each $\omega\in\omega(\mathcal{A})$ and $\sigma\in H_{K}$, we let $\omega^{\sigma}$ be the base change of $\omega$ viewed as an element of $\omega(\mathcal{A})\otimes_{\mathcal{O}_{K},\sigma}R\cong H^{0}(A^{\sigma},\Omega^{g}_{A^{\sigma}/R})$, where $R=\overline{\mathbb{Q}_{p}}$ or $\mathbb{C}$ and $A^{\sigma}\coloneqq A\otimes_{K,\sigma}R$. We view $\omega(\mathcal{A})$ as a lattice in $H^{0}(A,\Omega_{A/K}^{g})$. This lattice allows us to define a valuation on $H^{0}(A,\Omega_{A/K}^{g})$ for each rational prime $p$ and embedding $\sigma\in\operatorname{Hom}(K,\overline{\mathbb{Q}_{p}})$ by setting $v_{p}(\omega^{\sigma})=0$ if $\omega^{\sigma}$ generates $\omega(\mathcal{A})\otimes_{\mathcal{O}_{K},\sigma}\mathcal{O}_{p}$ as an $\mathcal{O}_{p}$-module and $v_{p}(\alpha\omega^{\sigma})=v_{p}(\sigma(\alpha))+v_{p}(\omega^{\sigma})$ for any $\alpha\in K$. This gives us the local contribution to the height.
We now define a norm for each archimedean place $\sigma\colon K\to\mathbb{C}$. We can view $\omega^{\sigma}\in H^{0}(A^{\sigma},\Omega_{A^{\sigma}/\mathbb{C}}^{g})$ as a holomorphic $g$-form on $A^{\sigma}$ and we set
@@
[…]
Remark.
Our choice of norm at infinity differs from that of [YZ18]. Their metric, which we denote by $\lVert\omega^{\sigma}\rVert_{YZ}$, is given by
$\lVert\omega^{\sigma}\rVert_{YZ}=\frac{1}{(2\pi)^{g/2}}\lVert\omega^{\sigma}\rVert.$
Thus, the Faltings height given in [YZ18], which we denote by $h_{YZ}$, differs from the one used in this article in the following way:
$h_{YZ}(A)=h(A)-\frac{g}{2}\log 2\pi.$